x⁴+x²y²+y⁴ = 133
x²+xy+y² = 19

Solve for x and y
1
Expert's answer
2020-07-12T17:37:53-0400

Let "x^2+y^2=u, xy=v." Then "x^4+x^2y^2+y^4=(x^2+y^2)^2-x^2y^2=u^2-v^2"


"u^2-v^2=133""u+v=19"

"(u-v)(u+v)=133""u+v=19"

"u-v=7""u+v=19"

"u=13""v=6"

Hence


"x^2+y^2=13""xy=6"

"(x+y)^2 -2xy=13""xy=6"

"(x+y)^2=25""xy=6"

If "x+y=5"


"(2, 3), (3, 2)"

If "x+y=-5"


"(-3, -2), (-2, -3)"

Therefore


"(-3, -2), (-2, -3), (2, 3), (3, 2)"